OpenStudy (anonymous):

If a tennis ball has a diameter of approximately 6.9 centimeters, what is its surface area?

OpenStudy (katrinakaif):

How do we calculate the SA of a Sphere?

OpenStudy (katrinakaif):

The formula?

OpenStudy (bahrom7893):

TENNIS!

OpenStudy (bahrom7893):

SA = 4*pi*r^2 r=d/2 find SA

OpenStudy (anonymous):

598.28 square centimeters